OpenStudy (anonymous):

wondering if these equations will form a linear system and if yes will there be no solution or one or many solutions? 3z + x = -4 y + 5z = 1 6x + 2z = 3 -x -y -z = 4

OpenStudy (anonymous):

in linear algebra it should look like \[\left[\begin{matrix}1 & 0 & 3 \\ 0 & 1 & 5 \\ 6 & 0 & 2 \\ -1 & -1 & -1 \end{matrix}\right]\cdot \left(\begin{matrix}x \\ y \\ z \end{matrix}\right) = \left(\begin{matrix}-4 \\ 1 \\ 3 \\ 4 \end{matrix}\right)\] you have to find out if the matrix has full rank. if so look in a linear algebra book about solvability. (tip: gaussian algorithm)

OpenStudy (anonymous):

mmm mind if you can explain when will equations form a linear system?

OpenStudy (anonymous):

hmm, i don't really know what to say. "when its linear" but this wouldn't help you. a funktion \(f\) is linear if \[ f (a + b) = f(a) + f(b) \]and \[ f( a \cdot b ) = f(a) \cdot f(b) \] and a linear system is the funktion that solves the equations, and is linear. does that help you?

OpenStudy (anonymous):

mm well one thing im trying to understand is can a set of equations have difference variables or it has to have same variables to become a linear sytem. for e.g will 2x - 4y = 2 2x - 4y + 5z = 5 be a linear system if the first one has 2 unknowns and second one has 3

OpenStudy (anonymous):

\[2x - 4y + 0z = 2 \]\[2x - 4y + 5z = 5 \]here its easy you can set z to \(3/5\)

OpenStudy (anonymous):

alright i see

OpenStudy (anonymous):

well thanks gonna post another questions :P

OpenStudy (anonymous):

you're welcome. please mark me as good. i just started with openstudy, i like it.

OpenStudy (anonymous):

what about x1 + x2 = x3 + x4? I think it will form a linear system and since it has one unknown then it will have infinite solutions right?

OpenStudy (anonymous):

sorry im wrong it has 4 unknowns

OpenStudy (anonymous):

yeah if you want to solve linear problems pretty easy. use matrices. like i did in my first post. it's pretty easy. but i must go now. i hope i could help you.